# Robust Inverse Graphics via Probabilistic Inference

### v1

@@ 7 Discussion
There are two key limitations and areas of improvement for our work. The reliance on a 3D prior makes it difficult to apply our method to robust perception of real world scenes, as it requires a 3D prior over real world scenes (see Appendix H for prior samples from models used in the experiments). While a lot of progress has been made in unconditional 3D models (Bautista et al., 2022; Chen et al., 2023; Chai et al., 2023; Kim et al., 2023), their fidelity and breadth of coverage of real world scenes remains small.
@@
Broader Impacts
@@ Appendix B ReGAL with Importance Sampling
ReGAL (Section 4, Algorithm 1) can be viewed as a one-sample importance sampler that samples from an extended-space target density
@@
We call this importance sampler ReGAL-IS and show the full algorithm in Algorithm 2.
@@ Appendix C ReGAL with Sequential Monte Carlo
Given the sequential nature of ReGAL and ReGAL-IS, it is natural to introduce ReGAL within a sequential Monte Carlo (SMC) framework (Doucet et al., 2009; Chopin et al., 2020) (ReGAL-SMC) which generalizes and improves upon the previous versions.
Both ReGAL and ReGAL-IS sample variables in the following order
@@
Given a sequence of variables $z^{(1:N)},c^{(1:N)}$, SMC sequentially builds approximations of target distributions $\gamma_{n}(z^{(1:n)},c^{(1:n)})$ from $n=1$ to $N$. Since the final target distribution $\gamma_{N}$ in (13) marginalizes to the desired posterior distribution $p(x,c|y)$ (like in ReGAL-IS), given an approximate sample $(z^{(1:N)},c^{(1:N)})$ from $\gamma_{N}$, we can take $(x,c):=(z^{(N)},c^{(N)})$ to be the approximate sample from $p(x,c|y)$.
@@ Appendix F Scene Prior Evaluation Details
We use ReGAL-IS for ShapeNet exclusively, and ReGAL-SMC for MSN. We use $K=8$ in all cases except clean ShapeNet where this is unnecessary due to the simplicity of the dataset. In general, simpler datasets/corruptions require fewer particles as the posterior is simpler.
Appendix G PSNR Evaluation
Figure 7: PSNR histograms across all conditions. Higher is better.
ShapeNet
Model Clean Rain Cloud FOV
ProbNeRF MAP $27.00\pm 0.37$ $19.00\pm 0.68$ $17.00\pm 0.51$ $32.00\pm 0.85$
ProbNeRF VI $24.00\pm 0.79$ $23.00\pm 0.64$ $19.00\pm 0.61$ $33.00\pm 1.0$
SSDNeRF ReGAL $\mathbf{31.00\pm 0.36}$ $\mathbf{26.00\pm 0.55}$ $\mathbf{20.00\pm 0.61}$ $\mathbf{35.00\pm 0.57}$
Table 2: Mean PSNR values across all conditions. Higher is better. The confidence intervals are 3 $\times$ SEM computed from a single trial via bootstrap. Note that the PSNR values for the FOV dataset are inflated because of the white background which takes up a higher fraction of the image for larger fields-of-view values.
Appendix H Prior samples

### v2

@@ 7 Discussion
There are two key limitations and areas of improvement for our work. The reliance on a 3D prior makes it difficult to apply our method to robust perception of real world scenes, as it requires a 3D prior over real world scenes (see Appendix G for prior samples from models used in the experiments). While a lot of progress has been made in unconditional 3D models (Bautista et al., 2022; Chen et al., 2023; Chai et al., 2023; Kim et al., 2023), their fidelity and breadth of coverage of real world scenes remains small.
@@
Acknowledgements
Thanks to Jyh-Jing Hwang for his helpful comments and Mehdi S. M. Sajjadi for his helpful comments and help with setting up the scene representation transformer model.
Impact Statement
@@ Appendix B ReGAL with Importance Sampling
The basic variant of ReGAL (Section 4, Algorithm 1) can be viewed as a one-sample importance sampler that samples from an extended-space target density
@@
The full algorithm can be seen in Algorithm 2.
@@ Appendix C ReGAL with Sequential Monte Carlo
Given the sequential nature of ReGAL, it is natural to introduce place it within a sequential Monte Carlo (SMC) framework (Doucet et al., 2009; Chopin et al., 2020). The resultant algorithm, dubbed ReGAL-SMC, generalizes the previous versions.
ReGAL samples variables in the following order
@@
Given a sequence of variables $z^{(1:N)},c^{(1:N)}$, SMC sequentially builds approximations of target distributions $\gamma_{n}(z^{(1:n)},c^{(1:n)})$ from $n=1$ to $N$. Since the final target distribution $\gamma_{N}$ in (13) marginalizes to the desired posterior distribution $p(x,c|y)$, given an approximate sample $(z^{(1:N)},c^{(1:N)})$ from $\gamma_{N}$, we can take $(x,c):=(z^{(N)},c^{(N)})$ to be the approximate sample from $p(x,c|y)$.
@@ Appendix F Scene Prior Evaluation Details
Appendix G Prior Samples
@@
Appendix H Additional Reconstructions
See Figures 11, 12, 13 and 14.
Figure 11: Reconstructions for the rain corruption of the ShapeNet cars dataset. See Figure 4 for details.
Figure 12: Reconstructions for the clouds corruption of the ShapeNet cars dataset. See Figure 4 for details.
Figure 13: Reconstructions for the rain corruption of the MultiShapeNet dataset. See Figure 4 for details.
Figure 14: Reconstructions for the clouds corruption of the MultiShapeNet. See Figure 4 for details.
Appendix I Posterior Uncertainty
In Figure 15 we show multiple samples from the posterior for the clouds corruption of the ShapeNet cars dataset for SSDNeRF ReGAL ($K=8$).
Figure 15: Multiple samples from the posterior for the clouds corruption of the ShapeNet cars dataset for SSDNeRF ReGAL (K=8). See Figure 4 for details.
Appendix J Performance
On a single A100 GPU, for each image it takes 9.5 minutes for MultiShapeNet and 7.5 minutes for ShapeNet to run ReGAL for 2000 steps to generate 8 particles. The time is proportional to the number of steps, so at a cost of quality, it could be reduced (as much as 10x before complete breakdown). The particles can be simulated on separate GPUs as well with relatively small cross-GPU communication requirements (due to IS/SMC resampling). Ultimately, however, practical deployment of this method would require more advanced diffusion sampling methods (such as Consistency Models (Song et al., 2023) and Flow Matching (Lipman et al., 2022)) which would drastically reduce the number of steps we need to take. It is likely these are not drop-in and would require adjustment of the ReGAL algorithm.